OpenStudy (anonymous):

How do I simplify: (x^2-6x+8/2x^2-7x+6) - (x^2-x-12/2x^2-5x-12)?

OpenStudy (anonymous):

x^2-6x+8 x^2-x-12 --------- - ---------- 2x^2-7x+6 2x^2-5x-12 = (x-4)(x-2) (x-4)(x+3) ---------- - --------- (2x-3)(x-2) (2x+3)(x-4) = (x-4)(2x+3) (x-4)(x+3) ----------- - ---------- (2x-3)(2x+3) (2x+3)(2x-3) = (x-4)(2x+3) - (x+3)(2x-3) ------------------------- (2x-3)(2x+3) = 2x^2-5-12- (2x^2+3x-9) --------------------------- (2x-3)(2x+3) = -5x-12-3x+9 ------------- (2x-3)(2x+3) = -8x-3 ----------- (2x-3)(2x+3) Is this right?

OpenStudy (turingtest):

No, let me write it x^2-6x+8 x^2-x-12 --------- - ---------- (1) 2x^2-7x+6 2x^2-5x-12 = (x-4)(x-2) (x-4)(x+3) ---------- - --------- (2) (2x-3)(x-2) (2x+3)(x-4) so far so good, but now we should notice that x-2 cancels in this first expression... (continued)

OpenStudy (turingtest):

and the x+4 in the second, giving\[{x−4\over2x+3}−{x+3\over2x+3}={x−4−x−3\over2x+3}={−7\over2x+3}\]which is your answer

OpenStudy (anonymous):

ohhh...

OpenStudy (anonymous):

really?